In a scene from the movie 'Matrix Reloaded', Trini
Answer #1

this is the problem of doppler effect

relavent formula for this

fs=(C/C-Vs)*f0

fO is the observed frequency

fS is the emitted frequency

c is the velocity of sound =340 m/s

714 = (340/(340-V)*710

V= 1.905 m/s

thus speed of agent = V+54 =55.905 m/s

=